on graph paper draw the graph of y = 2 sin(pi times x) for 0, or = x , or = 2 times pi and identify amplitude, period and the number of cycles completed in the given interval??? I have no idea how to even begin, Any assistance is appreciated :)
Answer by lwsshak3(6460)

on graph paper draw the graph of y = 2 sin(pi times x) for 0, or = x , or = 2 times pi and identify amplitude, period and the number of cycles completed in the given interval
**
Equation for graphing sin function: y=Asin(Bx-C), A=amplitude, Period=2π/B, Phase shift=C/B.
For given equation, y=2sin(πx) for 0 ≤ x ≤ 2π
Amplitude=2
B=π
Period=2π/B=2π/π=2
1/4 Period=2/4=1/2
C=0
Phase shift: none
Graphing:
On an (x,y) coordinate system with the x-axis labeled in radians, make the following tick marks on the x-axis: 1/2, 1, 3/2, 2, 5/2, 3, 7/2, 4 (this covers 2 periods)
Plot the following points: (0,0), (1/2,2), (1,0), (3/2,-2), (2,0), (5/2,2), (3,0), (7/2,-2), (4,0)
I don't have the means to graph it for you but you have the info you need to graph given sin function. Note: Your graph should have the same shape as the basic sin curve.